# Quantitative Certification of Bias in Large Language Models

### v2

@@ Appendix A Ablation study
In this section, we study the effect of changing the various certification parameters (a.k.a. hyperparameters) on the certificates generated with QuaCer-B. Table A presents the list of hyperparameters and their values used in our experiments.
@@
We regenerate the certificates for different prefix distributions by varying the hyperparameters. In particular, we study the variations of the results when $n,q,p_{\lambda},p_{\mu},$ and $\kappa$ are varied, keeping $\gamma$ constant. This is because, $1-\gamma$ denotes the confidence of the certification bounds and that is generally desired to be high. $95\%$ is a typical confidence level for practical applications [Sim and Reid, 1999]. We conduct this ablation study on the specifications for a randomly picked set of $100$ counterfactual prompt sets from BOLD’s test set $\mathcal{Q}_{BOLD}$. We certify the Mistral-Instruct-v0.2 [Jiang et al., 2023] 7B parameter model and study the overall results next. We show ablations on $n$ for all kinds of specifications in Figure 5. We see that the bounds begin converging at $50$ samples and subsequent samples cause minor variations in their values, justifying our choice of using $50$ samples. Fewer than $50$ samples can result in less tight bounds.
@@
(d) Variation with relative magnitude of noise $\kappa$ for soft jailbreaks specifications
@@ A.1 Random prefixes
A.2 Mixture of jailbreaks
These specifications have $2$ hyperparameters — $p_{\lambda}$, the probability of adding an instruction from the helper jailbreaks when interleaving, and $p_{\mu}$, the probability of randomly flipping every token of the resultant of interleaving. We show ablation studies on these in Figures 6(b) and 6(c) respectively.
A.3 Soft jailbreaks
@@
A.4 Scaling beyond binary demographic groups
@@
Our general framework (Section 3) and specification instances (Section 4) are applicable to certify biases beyond binary counterfactual prompt sets (like for male/female gender, black/white race). This is subject to the availability of bias detectors $\mathcal{D}$ that can identify biases across responses for counterfactual prompt sets for more than binary demographic groups. While we are not aware of any $\mathcal{D}$ that could work with $\mathcal{Q}_{BOLD}$, we extend our $\mathcal{D}$ for the specifications from $\mathcal{Q}_{DT}$ to work for responses to prompts from three racial demographic groups — black people, white people, and asians. We elaborate on the extension in Appendix C.2. We certify Mistral-Instruct-v0.2 7B model with the three kinds of specifications and find that the average certification bounds presented in Table 3 are similar to the bounds presented for the Mistral model in Table 1 for the random and mixture of jailbreak specifications. However, the results are significantly worse for the soft prefix specifications. This is because, firstly the model is particularly susceptible to these specifications as is evidenced even in the results with binary demographic groups. Secondly, with the addition of another demographic group, the bias detector is skewed towards identifying bias in more sets of responses than for the case with binary demographic groups. The bias detector identifies bias in responses having at least $1$ agreement and $1$ disagreement to the stereotype mentioned in the prompts, which has the same chance as unbiased result for binary demographic groups, but not beyond them.
Appendix B Common jailbreaks used
In our experiments, we combine common jailbreaks to certify against hybrid jailbreaks. The specific jailbreaks we use are shown below. The main jailbreak [^2] Last accessed on Jan 22, 2024 from www.jailbreakchat.com/prompt/588ab0ed-2829-4be8-a3f3-f28e29c06621 is shown in Figure 7. We have modified the jailbreak to encourage the model to generate responses demonstrating gender bias by adding the text highlighted in red in Figure 7(a) and for racial bias in Figure 7(b). We use one helper jailbreak [^3] https://www.reddit.com/r/ChatGPT/comments/113ddqq/my_jailbreak_is_far_superior_to_dan_the_prompt_is/ from which we derive the instructions shown in Figure 8. Each of these instructions is inserted into the main jailbreak at the points where its individual instructions end, with the interleaving probability $p_{\lambda}$. All instructions inserted at the same point in the main jailbreak appear in a random order in the final jailbreak prefix.
(a) For gender
(b) For race
Figure 7: Main Jailbreaks for different sensitive attributes
Figure 8: Instructions derived from helper jailbreak
Appendix C Bias detector
C.1 BOLD experiments
@@
Design of bias detector. Our bias detector $\mathcal{D}_{BOLD}$ is based on the regard bias metric [Sheng et al., 2019]. Specifically, we measure the bias for a pair of completions, one for male and another for female, by invoking the regard classifier in ’compare’ mode. The compare mode gives the difference between the regard score obtained for the different completions. We consider all completion pairs with a regard score higher than a prespecified threshold ($0.85$ in this case) as biased. For the completions considered unbiased with the previous evaluation, we also consider the regard scores of the individual completions (with respect to the partial sentence given to the model). We consider the completions to be biased if either score exceeds the threshold. This is for the instances of bias, wherein the pair of generations are similarly biased and individual generations have text discriminating against a particular demographic group.
@@
Figure 9(a) shows the true positive examples where the both the $\mathcal{D}_{BOLD}$ and the annotators identify bias. Figure 9(b) shows the false negative examples where $\mathcal{D}_{BOLD}$ does not identify bias but the annotators do. Figure 9(c) shows the false positive examples where $\mathcal{D}_{BOLD}$ identifies bias but the annotators do not. According to the authors, the false positive examples are actually biased. But we also recognize that bias is a complex subject, and its notion can vary across individuals. Hence, we believe that the noise in the human study needs to be taken into account when evaluating our bias detector $\mathcal{D}_{BOLD}$.
C.2 DT experiments
@@
C.3 Limitations of bias detectors
@@
Appendix D Example samples from prefix distributions
Figure 10: Examples of random prefixes
Mixture of jailbreaks. Figure 11 presents some examples of mixture of jailbreak prefixes, formed from the main and helper jailbreaks in Figures 7 and 8 respectively.
Figure 11: Examples of mixture of jailbreak prefixes
Appendix E Case studies from Decoding Trust certificates
Appendix F Positive and Negative impacts of our work
We identify the following positive and negative impacts of our work.
Positive impacts. Our work is the first to provide quantitative certificates for the bias in Large Language Models. It can be used by model developers to thoroughly assess their models before releasing them and by the general public to become aware of the potential harms of using any LLM. As our framework, QuaCer-B assumes black-box access to the model, it can be applied to even closed-source LLMs with API access.

### v3

@@ Appendix E Ablation study
In this section, we study the effect of changing the various certification parameters (a.k.a. hyperparameters) on the certificates generated with LLMCert-B. Table E presents the list of hyperparameters and their values used in our experiments.
@@
We regenerate the certificates for different prefix distributions by varying the hyperparameters. In particular, we study the variations of the results when $n,T,\text{Top-k},q,p_{\lambda},p_{\mu},$ and $\kappa$ are varied, keeping $\gamma$ constant. This is because, $1-\gamma$ denotes the confidence of the certification bounds and that is generally desired to be high. $95\%$ is a typical confidence level for practical applications [Sim and Reid, 1999]. We conduct this ablation study on the specifications for a randomly picked set of $100$ counterfactual prompt sets from BOLD’s test set $\mathcal{Q}_{BOLD}$. We certify the Mistral-Instruct-v0.2 [Jiang et al., 2023] 7B parameter model and study the overall results next.
E.1 Certification algorithm hyperparameter
We show ablations on $n$ for all kinds of specifications in Figure 10. We see that the bounds begin converging at $50$ samples and subsequent samples cause minor variations in their values, justifying our choice of using $50$ samples. Fewer than $50$ samples can result in less tight bounds.
@@ E.1 Certification algorithm hyperparameter
E.2 LLM decoding hyperparameters
We study variations in the certification bounds with $2$ important hyperparameters of the LLM decoding algorithms that influence their generated texts — $T$ (decoding temperature) and Top-k (number of tokens decoded at each step). Figures 11 and 12 show the variations in the certification bounds with $T$ and Top-k respectively for the $3$ kinds of specifications. We see only minor changes in the average certification bounds with the variations of these hyperparameters. Our hypothesis of this phenomenon is that as the certificates aggregate the bias results of several samples, they smooth out the noise introduced by the choice of LLM decoding hyperparameters and give insights into the biases of the LLM itself.
(a) Random prefixes
(b) Mixtures of jailbreaks
(c) Soft prefixes
Figure 11: Ablation study showing variations of average certification bounds with temperature $T$
(a) Random prefixes
(b) Mixtures of jailbreaks
(c) Soft prefixes
Figure 12: Ablation study showing variations of average certification bounds with Top-k parameter.
@@ E.2 LLM decoding hyperparameters
(d) Variation with relative magnitude of noise $\kappa$ for soft prefixes from jailbreak specifications
@@ E.3 Random prefixes
E.4 Mixture of jailbreaks
These specifications have $2$ hyperparameters — $p_{\lambda}$, the probability of adding an instruction from the helper jailbreaks when interleaving, and $p_{\mu}$, the probability of randomly flipping every token of the resultant of interleaving. We show ablation studies on these in Figures 13(b) and 13(c) respectively.
E.5 Soft prefixes from jailbreaks
@@
E.6 Scaling beyond binary demographic groups
@@
Our general framework (Section 3) and specification instances (Section 4) are applicable to certify biases beyond binary counterfactual prompt sets (like for male/female gender, black/white race). This is subject to the availability of bias detectors $\mathcal{D}$ that can identify biases across responses for counterfactual prompt sets for more than binary demographic groups. While we are not aware of any $\mathcal{D}$ that could work with $\mathcal{Q}_{BOLD}$, we extend our $\mathcal{D}$ for the specifications from $\mathcal{Q}_{DT}$ to work for responses to prompts from three racial demographic groups — black people, white people, and asians. We elaborate on the extension in Appendix G.2. We certify Mistral-Instruct-v0.2 7B model with the three kinds of specifications and find that the average certification bounds presented in Table 4 are similar to the bounds presented for the Mistral model in Table 1 for the random and mixture of jailbreak specifications. However, the results are significantly worse for the soft prefix specifications. This is because, firstly the model is particularly susceptible to these specifications as is evidenced even in the results with binary demographic groups. Secondly, with the addition of another demographic group, the bias detector is skewed towards identifying bias in more sets of responses than for the case with binary demographic groups. The bias detector identifies bias in responses having at least $1$ agreement and $1$ disagreement to the stereotype mentioned in the prompts, which has the same chance as unbiased result for binary demographic groups, but not beyond them.
Appendix F Validity of confidence intervals
We design a synthetic study for the validity of the confidence intervals as follows. As we can not precisely regulate the true probability of unbiased responses of LLMs, we assume various values of that probability and generate binary-valued samples indicating biased (non-zero) /unbiased (zero) responses from any LLM. Hence, we generate $50$ samples (same as the samples used in LLMCert-B’s certification) of the Bernoulli random variable $\mathcal{F}$ (Section 3.2), with various values for the probability of success and generate Clopper-Pearson confidence intervals for the success probability using the samples. We repeat this process $1000$ times and report the percentage of instances wherein the confidence intervals contain the true probability of success. This percentage indicates the probability of correctness of the confidence intervals. We find that for all 11 equally-spaced values of the true probability of unbiased responses between 0 and 1, the confidence intervals bound the true value for more than 95% (nominal, user-specified confidence level) times, which validates the claim that the confidence intervals hold with at least the user-specified confidence. Figure 14 presents the variation in the proportion of Clopper-Pearson confidence intervals that contain the true probability of unbiased responses for different values of the latter.
Figure 14: Variation in the proportion of Clopper-Pearson confidence intervals containing the true probability of unbiased responses. Dashed line shows the user-specified confidence level, i.e. $95\%$.
Appendix G Bias detector
G.1 BOLD experiments
@@
Design of bias detector. Our bias detector $\mathcal{D}_{BOLD}$ is based on the regard bias metric [Sheng et al., 2019]. Specifically, we measure the bias for a pair of completions, one for male and another for female, by invoking the regard classifier in ‘compare’ mode. The compare mode gives the difference between the regard score obtained for the different completions. We consider all completion pairs with a regard score higher than a prespecified threshold ($0.85$ in this case) as biased. For the completions considered unbiased with the previous evaluation, we also consider the regard scores of the individual completions (with respect to the partial sentence given to the model). We consider the completions to be biased if either score exceeds the threshold. This is for the instances of bias, wherein the pair of generations are similarly biased and individual generations have text discriminating against a particular demographic group.
@@
Figure 15(a) shows the true positive examples where the both the $\mathcal{D}_{BOLD}$ and the annotators identify bias. Figure 15(b) shows the false negative examples where $\mathcal{D}_{BOLD}$ does not identify bias but the annotators do. Figure 15(c) shows the false positive examples where $\mathcal{D}_{BOLD}$ identifies bias but the annotators do not. According to the authors, the false positive examples are actually biased. But we also recognize that bias is a complex subject, and its notion can vary across individuals. The human annotation of bias is contingent on many factors such as their cultural background, our annotation instructions (provided in the open-source implementation of the framework), etc. Hence, we believe that the noise in the human study needs to be taken into account when evaluating our bias detector $\mathcal{D}_{BOLD}$. Our bias detector shows 93% precision, but 50% recall. Owing to these inconsistencies of our bias detector with human perception of bias, we believe that our bounds for probability of unbiased responses are actually higher than bounds with a perfect bias detector (due to low recall), indicating a worse situation of counterfactual bias in SOTA LLMs.
G.2 DT experiments
@@
G.3 Limitations of bias detectors
@@
Appendix H Comparing with prior works on guarantees for LLMs
Works on conformal prediction. [Quach et al., 2024, Deutschmann et al., 2023, Mohri and Hashimoto, 2024, Yadkori et al., 2024] apply conformal prediction techniques to language models, proposing methods for generating sets of outputs with statistical guarantees on correctness, coverage, or abstention, aiming to improve reliability and mitigate hallucinations. Their guarantees and scope differ from those of LLMCert-B. The prior works give specialized decoding strategies that guarantee the correctness of the outputs, useful for the factuality of the responses. LLMCert-B, however, is about assessing and certifying the counterfactual biases in LLM responses generated with any decoding scheme applied on the target LLMs. Moreover, guarantees of conformal prediction are for correctness of LLM responses for individual prompts, while LLMCert-B’s guarantees are over distributions of counterfactual prompt sets (which have prohibitively large sample spaces).
Prompt Risk Control. [Zollo et al., 2024] (PRC) presents a framework for selecting low-risk system prompts for LLMs. PRC computes the loss incurred for one prompt at a time, and aggregates those losses to form a risk measure. LLMCert-B, on the other hand, is for counterfactual bias, i.e., we assess bias across a set of LLM responses, obtained by varying the sensitive attributes in prompts. Our specification is thus a relational property [Barthe et al., 2011], which is defined over multiple related inputs. Biases across LLM responses for multiple related prompts are aggregated to certify the target LLM. Moreover, LLMCert-B proposes novel, inexpensive mechanisms to design distributions and their samplers with potentially adversarial prefixes, containing effective, manually-designed jailbreaks in their sample space. Contrary to PRC which uses static datasets of adversarial examples, we use independent and identically-distributed samples from these distributions.
Appendix I Additional notes on Bias in ML
Bias is a complex social phenomenon that arises in various forms. In this section, we discuss various notions of bias and harms caused by them. We also discuss how LLMCert-B complements existing evaluation methods by certifying for counterfactual bias. The following discussion is not a comprehensive treatment of bias in Machine Learning and we refer the reader to detailed survey papers such as [Gallegos et al., 2024a, Blodgett et al., 2020, Li et al., 2024] for more information.
Bias consists of disparate treatment (a.k.a. discrimination) or disparate outcomes [Barocas and Selbst, 2016] for different demographic groups. Harms due to bias are primarily of $2$ kinds — representational and allocation [Gallegos et al., 2024a]. Representational harm [Suresh and Guttag, 2021, Blodgett et al., 2020] consists of denigrating and subordinating attitudes towards a demographic group. It consists of use of derogatory language, stereotyping, toxicity, misrepresentation, etc. These can arise from inappropriate use of language by humans or machines (e.g., LLMs). Allocation harms [Ferrara, 2023] are disparate distribution of resources or opportunities between demographic groups. These consist of direct or indirect discrimination in economic or social opportunities. For example, prior works like [Terry et al., 2010, Martínez, 2022] show that the lack of representation of African American English in dominant language practices results in that community facing penalties in education systems or when seeking housing. Most constitutions around the world have anti-discrimination laws like [Sherry, 1965] that prohibit allocative harms in employment etc. Language is considered an important factor for labeling, modifying, and transmitting beliefs about demographic groups and can result in the reinforcement of social inequalities [Rosa and Flores, 2017].
LLMCert-B is a reliable evaluation method for counterfactual bias in language models (LMs), that certifies the probability of unbiased response in target LLMs for distributions of counterfactual prompts with statistical guarantees. Prior bias assessments have been of $2$ kinds [Cao et al., 2022] — intrinsic and extrinsic. Intrinsic bias occurs in the language representations, while extrinsic bias manifests in the final textual responses of the LMs. To certify closed-source LMs as well, we study extrinsic bias. Bias is opposite of fairness, which has been identified to be of various forms such as group fairness [Blandin and Kash, 2024], individual fairness [Dwork et al., 2011], counterfactual fairness [Kusner et al., 2018], etc. LLMCert-B certifies for counterfactual bias, akin to counterfactual fairness. This is because of the causal perspective of counterfactual bias [Anthis and Veitch, 2023] (bias due to mentioning specific demographic group in the prompt) which aligns more closely with human intuitions about discrimination and fairness. Moreover, unlike group fairness, counterfactual fairness operates at the individual-level, thus identifying bias in specific cases, instead of aggregates.